Determine, without graphing, whether this function has a maximum or minimum value. Then, find the value: f(x) = 3x2 + 5x – 7
Answer by 303795(602)

The a value in the equation (3) is positive so the equation will have a minimum. (A negative a value gives a maximum).
The minimum will be on the line of symmetry which is found using
LOS =-b/2a
so the x value at the minimum will be -5/(2*3)=-5/6.
Substitute x = -5/6 into the equation to get the y value of -(9 and 1/12)